You have 50 microwaves. Out of the 50 you have 4 defective ones. What is the probability that 2 out of the first 5 microwaves sold are defective? and how do I set up the equation?
Answer by stanbon(57361)

You have 50 microwaves. Out of the 50 you have 4 defective ones. What is the probability that 2 out of the first 5 microwaves sold are defective? and how do I set up the equation?
------------------
This is a binomial distribution problem.
P(defective) = 4/50
P(not defective) = 46/50
-------------------------------
P(2 of 5 are defective) = 5C2*(4/50)^2*(46/50)^3 = 0.049836...
Cheers,
Stan H.